OpenStudy (anonymous):

Given the sequence 12, 6, 3, 3/2, ..., identify the common ratio.

OpenStudy (alexandervonhumboldt2):

look: 12/6=2 6/3=2 3/2=3/2 any ideas?

OpenStudy (jcarangui1):

divide by two to get the next number

OpenStudy (alexandervonhumboldt2):

division of each number by the next equals 2, thus the common ration is?

OpenStudy (anonymous):

2 ?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

noon wait

OpenStudy (alexandervonhumboldt2):

@Starbucks wait

OpenStudy (alexandervonhumboldt2):

the ratio would be 1/2 as the sequence is `from biggest to less`

OpenStudy (anonymous):

ohhhh okay thank you